### v1

@@ Abstract
Recent advances in large language models (LLMs) provide a new opportunity to address the long-standing question of how concepts are represented and organized in the mind, which is central to unravelling the nature of human cognition. Here, we reframed the classic reverse dictionary task to simulate human concept inference in context and investigated the emergence of human-like conceptual representations within LLMs. We found that LLMs were able to infer concepts from definitional descriptions and construct representation spaces that converge towards a shared, context-independent structure. These representations effectively predicted human behavioural judgments and aligned well with neural activity patterns in the human brain, offering evidence for biological plausibility. These findings demonstrate that human-like conceptual representations and organization can naturally emerge from language prediction, even without real-world grounding. Our work supports the view that LLMs serve as valuable tools for understanding complex human cognition and paves the way for better alignment between artificial and human intelligence.
@@ 1 Introduction
Humans can construct mental models of the world and use them to understand and navigate their environment [1, 2]. Central to this ability is the capacity to form broad concepts that constitute the building blocks of these models [3]. Often regarded as mental representations, these concepts capture regularities in the world while abstracting away extraneous details, enabling flexible generalization to novel situations [4]. For example, the concept of sun can be instantly formed and deployed across diverse contexts: observing it rise or set in the sky, yearning for its warmth on a chilly winter day, or encountering someone who exudes positivity. Its role in the solar system can be analogized to the nucleus in an atom, enriching learning and understanding. The nature of concepts has long been a focus of inquiry across philosophy, cognitive science, neuroscience, and linguistics [5, 6, 7, 8, 9, 10, 11, 12]. These investigations have uncovered diverse properties that concepts need to satisfy, often framed by the long-standing divide between symbolism and connectionism. Symbolism emphasizes discrete, explicit symbols with structured and compositional properties, enabling abstract reasoning and recombination of ideas [13, 14]. In contrast, connectionism conceptualizes concepts as distributed, emergent patterns across networks, prioritizing continuity and gradedness, which excel in handling noisy inputs and learning from experience [15, 16]. Although there is growing consensus on the need to integrate the strengths of both paradigms to account for the complexity and richness of human concepts [17, 18, 19], reconciling the competing demands remains a significant challenge.
@@
Recent advances in large language models (LLMs) within artificial intelligence (AI) have exhibited human-like behaviour across various cognitive and linguistic tasks, from language generation [20, 21, 22] to decision-making [23] and reasoning [24, 25, 26, 27]. It has sparked intense interest and debate about whether these models are approaching human-like cognitive capacities based on concepts [28, 29, 30, 31, 32, 33]. Some argue that LLMs, trained exclusively on text data for next-token prediction, operate only on statistical associations between linguistic forms and lack concept-based understanding grounded in physical and social situations [34, 35, 32]. This argument is evidenced by their inconsistent performance, which often reveals vulnerabilities such as non-human-like errors and over-sensitivity to minor contextual variations [23, 36, 37]. Conversely, others contend that the performance of a system alone is insufficient to characterize its underlying competence [38], and the extent to which concepts should be grounded remains open [39, 40]. Instead, language may provide essential cues for people’s use of concepts and enable LLMs to approximate fundamental aspects of human cognition, where meaning arises from the relationships among concepts [41, 42]. Despite the conflicting views, there is broad consensus on the central role of concepts in human cognition. The core questions driving the debates are whether LLMs possess human-like conceptual representations and organization and, consequently, whether they can illuminate the nature of human concepts.
@@
To address these questions, we investigated the emergence of human-like conceptual representations from language prediction in LLMs. Our approach unfolded in three stages. First, we examined LLMs’ capacities to derive conceptual representations, focusing on the definitional and structured properties of human concepts. We leveraged LLMs’ in-context learning ability and guided them to infer concepts from descriptions with a few contextual demonstrations (Fig. 1). The models’ outputs offered a behavioural lens to evaluate the derived representations. We explored the organization of these representations by analysing how their relational structures varied across different contextual demonstrations. Second, we assessed how well the LLM-derived representational structures align with psychological measures and investigated whether they capture rich, context-dependent human knowledge. These questions were tested by leveraging computations over the representations to predict human behavioural judgments. Finally, we mapped the LLM-derived conceptual representations to neural activity patterns in the human brain and analysed their biological plausibility. Our experiments spanned thousands of naturalistic object concepts and beyond [43], providing a comprehensive analysis of LLMs’ conceptual representations. The findings reveal that language prediction can give rise to human-like conceptual representations and organization in LLMs. These representations integrated many valued aspects of previous accounts, combining the definitional and structural focus of symbolism with the continuity and graded nature of connectionist models. This work underscores the profound connection between language and human conceptual structures and demonstrates that LLM-derived conceptual representations offer a promising foundation for understanding human concepts.
@@ 2.1 Reverse dictionary as a conceptual probe
To guide LLMs through the process, we took advantage of their in-context learning ability and presented them with a small number of demonstrations in a reverse-dictionary format (“$[\textrm{Description}]\Rightarrow[\textrm{Word}]$”), followed by a query description (Fig. 1). We then compared model-generated completions to the intended term of the concept corresponding to the query description, thereby evaluating how well the models derived conceptual representations that aligned with human understanding. Formally, given $N$ pairs of descriptions and words as contextual demonstrations, an LLM $\mathcal{M}$ encodes the query description of a concept $c$ into a representation $\mathbf{h}_{c}=\textrm{encode}_{\mathcal{M}}\left(\textrm{description}\mid\textrm{context}_{N}\right)$, and generates a term based on it: $\textrm{predict}_{\mathcal{M}}\left(\mathbf{h}_{c}\right)$. The term was then compared to the name of the query concept shared by humans: $\textrm{name}_{\mathcal{H}}\left(c\right)$. We took $\mathbf{h}_{c}$ as the representation for concept $c$, which immediately precedes the following term that should be in semantic correspondence to the query description. Importantly, the demonstrations provided minimal and controllable context, and the query description served as a special case of general input for concept inference. We can analyse how models construct conceptual representations in response to different contextual cues by varying the number of demonstrations and the specific description-word pairings within them.
@@ 2.2 Deriving concepts from definitional descriptions
As a follow-up, we conducted a counterfactual analysis to investigate how LLMs accomplish this: Do they merely function as lookup tables, rigidly mapping descriptions to words, or do they contextually infer concepts based on their interrelationships? As shown in Fig. 2, when presented with only a single demonstration identical to the query description, the model replicated the context, even though the description was paired with a proxy symbol instead of the correct term. As additional correct demonstrations of other concepts were introduced, the model gradually shifted from replication to predicting the proper word for the query concept. This behaviour persisted across various proxy symbols, suggesting that contextual cues about other concepts successfully guided the model in disregarding misleading information. When the correct demonstration for the query concept was provided, model performance slightly declined as more demonstrations were added, eventually dropping below the average in the standard generalization setting. This decline indicated that subtle conflicts among demonstrations could arise, with the influence of other concepts overshadowing that of the identical query concept. Collectively, these findings underscore the intricate interrelationships among concepts and their pivotal role in shaping model inference.
Figure 2: Performance of LLaMA3-70B on the reverse dictionary task measured through exact match accuracy. Across the four plots, the black lines denote the model’s performance when presented with $N$ demonstrations sampled from the training set and evaluated on an independent test set. The blue and red lines illustrate the model’s behaviour when presented with one misleading demonstration—a description paired with a proxy label (as specified in the figure titles)—and $N-1$ correct demonstrations of other concepts. The blue line shows the frequency with which the model copies the misleading demonstration and reproduces the proxy label, while the red line indicates how often it generates the correct word for the query concept given contextual information from other concepts. Shaded areas denote 95% bootstrapped confidence intervals, calculated from 10,000 resamples over five independent runs.
@@ 2.3 Uncovering a context-independent conceptual structure
To uncover how concepts are represented and organized within LLMs, we looked into the representation spaces they constructed for concept inference and analysed the interrelationships among the conceptual representations. We characterized the representation spaces formed under different contexts by their relational structure, captured through a (dis)similarity matrix, and measured pairwise alignment by correlating these matrices [46, 47] (Methods 4.3). For concepts in the THINGS database, the alignment between representation spaces gradually increased as the number of demonstrations rose from 1 to 24, with diminishing gains beyond this threshold (Fig. 3a). The alignment with the space formed by 120 demonstrations increased from $0.800$ (±$\pm 0.017$) with one demonstration to $0.970$ (±$\pm 0.003$) after 24 demonstrations. This alignment demonstrated a strong correlation with the model’s exact match accuracy on concept inference ($\rho=0.976$, $P<0.0001$, $95\%$ CI: 0.7300.7300.7300.7301.01.01.01.0) (Fig. 3b). These results suggest that LLMs are able to construct a coherent and context-independent relational structure, which is reflected in their concept inference capacity. Additional metrics corroborated this conclusion (Methods 4.3 and Supplementary Information SI 1.2), highlighting that intricate relationships among concepts are consistently preserved within LLMs’ representation spaces across varying contexts. This invariance supports the generalization of knowledge encoded in these relationships.
@@
Figure 3: LLMs converge toward a similar representational structure of concepts. a, Alignment correlation (RSA) between the LLM-derived conceptual representations across different contextual demonstrations. The LLM (LLaMA3-70B) is presented with $N$ demonstrations, repeated across five independent runs. b, LLM performance on the reverse dictionary task reflects alignment with the representations formed based on 120 demonstrations. Each point corresponds to the representations formed by the LLM based on $N$ demonstrations, with the x-axis indicating exact match accuracy and the y-axis denoting the alignment correlation. Error bars indicate $95\%$ confidence intervals, calculated using 10,000 bootstrap resamples from five independent runs. c, A t-SNE visualization of conceptual representations produced by different LLMs. Each point corresponds to an LLM, with representations formed based on 24 demonstrations. Distances between them are calculated as $1-\textrm{alignment}$, with results averaged over five independent runs. LLMs are plotted in proportion to their complexity and colour-coded by their exact match accuracy on the reverse dictionary task. Better-performing models (blue) exhibit more aligned representations. d, LLMs’ concept inference capacity reflects alignment correlation with conceptual representations derived from LLaMA3-70B. The x-axis represents the average performance on the reverse dictionary task with 24 demonstrations, while the y-axis indicates the alignment correlation with LLaMA3-70B representations. Data points correspond to conceptual representations derived from different models and are colour-coded by model complexity. Linear fits are shown as straight lines, with shaded areas representing $95\%$ confidence intervals derived from 10,000 bootstrap resamples. Overall, models with higher complexity align more closely with the reference. An exception is Qwen2-0.5B, which, despite exhibiting relatively high complexity, has a low alignment score. This may be due to constraints in model scale and the quality of training data.
@@ 2.4 Predicting various facets of human concept usage
Next, we investigated how well the conceptual representations and structures derived from LLMs align with various aspects of human concept usage. Specifically, we used these representations to predict human behavioural data across three key psychological phenomena: similarity judgments, categorization and gradient scales along various features. To evaluate their effectiveness, we compared the LLM-derived representations against traditional static word embeddings learned from word co-occurrence patterns [49].
@@
For similarity judgments, we compared LLM-derived conceptual representations with human similarity ratings for concept pairs from SimLex-999 [50], which has proven challenging for traditional word embeddings. We also complemented our analysis using human triplet odd-one-out judgments from the THINGS-data collection [43], which relied on image-based rather than text-based stimuli and introduced a contextual effect through the third concept. As shown in Fig. 4a, similarity scores derived from LLM representations strongly correlated with human ratings in SimLex-999, with the correlation improving as the number of contextual demonstrations increased ($\rho=0.776\pm 0.007$ with 72 demonstrations across five runs, $P<0.0001$). These representations significantly outperformed traditional word embeddings, which achieved a correlation of $\rho=0.464$ ($P<0.0001$). For THINGS triplets, we calculated pairwise similarities to identify the odd-one-out (Methods 4.5). The LLM prediction accuracy also improved with increasing contextual demonstrations, plateauing at $63.20\%$ (±$\pm 0.29\%$) after 48 demonstrations (Fig. 4b). This performance closely approached the noise ceiling estimated from individual behaviour ($67.67\%\pm 1.08\%$) and substantially exceeded that of word embeddings ($48.10\%$). These findings suggest that, within proper context, the conceptual representations formed by LLMs effectively support the computation of similarities, a core property of human concepts.
@@
We then examined whether categories can be induced from the relative similarity between LLM-derived conceptual representations, using the human-labelled high-level categories from the THINGS database [45]. Applying a prototype-based categorization approach (Methods 4.6), we observed that LLM-derived representations consistently achieved high accuracy, reaching $92.25\%$ (±$\pm 0.15\%$) with only 24 demonstrations, significantly outperforming static word embeddings ($77.88\%$) (see Supplementary Information SI 1.4 and Fig. S3c). A t-SNE visualization of these representations revealed notable differences in their similarity structures (Fig. 4c–d). The LLM-derived conceptual representations formed distinct clusters corresponding to high-level categories, such as animals and food, while word embeddings exhibited less distinct category separation. Furthermore, LLM representations revealed broader distinctions, separating natural and animate concepts from man-made and inanimate ones. Among these groupings, human body parts clustered more closely with man-made objects than other animals do, while processed food appeared closer to natural and animate objects. These patterns align with previous findings on human mental representations [51] and are prominent in LLM representations, underscoring the meaningful correspondence between LLM-derived conceptual structures and human knowledge.
@@
Figure 4: Evaluation of alignment between LLM-derived conceptual representations and psychological measures of similarity. ab, Performance of LLaMA3-70B conceptual representations compared to static word embeddings (Word) in predicting human similarity judgments. a, Performance on human similarity ratings for concept pairs from SimLex-999, evaluated by Spearman’s rank correlation. b, Performance on human triplet odd-one-out similarity judgments in THINGS, measured by prediction accuracy. The noise ceiling reflects the upper bound of performance based on individual response consistency for the same triplet. Error bars and shaded regions represent 95% confidence intervals calculated from five independent runs. cd, t-SNE plots of LLM-derived conceptual representations (c) and static word embeddings (d). Data points are colour-coded based on human-labelled categories from THINGS. Conceptual representations derived from LLaMA3-70B show a clear alignment with category structure, in contrast to the less distinct clustering observed with word embeddings.
@@
Finally, we investigated whether LLM conceptual representations could capture gradient scales of concepts along various features. For example, on a scale of 1 to 5 relative to other animals, cheetahs might rank a 5 for speed but fall closer to 3 for size. Using LLM-derived representations for such ratings, we predicted human behavioural data across 52 category-feature pairs (e.g., animals rated for speed) [52] (Methods 4.7). The results (Fig. 5) revealed strong correlations with human ratings ($\rho>0.5$, $P<0.001$, FDR controlled) for 48 out of 52 category-feature pairs, with a median correlation of $0.817$ ($95\%$ CI: 0.7770.7770.7770.7770.8510.8510.8510.851) across all pairs. When accounting for the split-half reliability of human ratings (median $\rho=0.954$), the median correlation reached $0.871$ ($95\%$ CI: 0.8410.8410.8410.8410.8910.8910.8910.891). Among the three category-feature pairs without statistically significant correlations (FDR $P>0.01$), two exhibited marginal significance (FDR $P<0.05$), while the weakest correlation appeared in the rating of clothing by age. As shown in Fig. 6, the LLM conceptual representations outperformed static word embeddings [52] across most category-feature pairs. This advantage remained robust after excluding concepts with extreme feature values from the correlation analysis (Supplementary Information SI 1.5), confirming that the success of LLMs is not driven by outlier items. Our findings demonstrate that LLM-derived conceptual representations advantageously handle intricate human knowledge across diverse object categories and features, highlighting their potential to advance our understanding of conceptual representations in the human mind.
@@
Figure 6: Comparison of LLM-derived conceptual representations and static word embeddings in predicting context-dependent human ratings. The x-axis represents correlations with human ratings based on static word embeddings, while the y-axis represents correlations based on conceptual representations derived from LLaMA3-70B. Coloured points indicate significant correlations (Spearman’s rank correlation, FDR $P<0.01$). Error bars represent 95% confidence intervals estimated from 10,000 bootstrap resamples.

### v3

@@ Abstract
People acquire concepts through rich physical and social experiences and use them to understand the world. In contrast, large language models (LLMs), trained exclusively through next-token prediction over language data, exhibit remarkably human-like behaviors. Are these models developing concepts akin to humans, and if so, how are such concepts represented and organized? To address these questions, we reframed the classic reverse dictionary task to simulate human concept inference in context and investigated the emergence of human-like conceptual representations within LLMs. Our results demonstrate that LLMs can flexibly derive concepts from linguistic descriptions in relation to contextual cues about other concepts. The derived representations converged towards a shared, context-independent structure that effectively predicted human behavior across key psychological phenomena, including computation of similarities, categories and semantic scales. Moreover, these representations aligned well with neural activity patterns in the human brain, even in response to visual rather than linguistic stimuli, providing evidence for biological plausibility. These findings establish that structured, human-like conceptual representations can naturally emerge from language prediction without real-world grounding. More broadly, our work positions LLMs as promising computational tools for understanding complex human cognition and paves the way for better alignment between artificial and human intelligence.
@@ 1 Introduction
Humans are able to construct mental models of the world and use them to understand and navigate their environment [1, 2]. Central to this ability is the capacity to form broad concepts that constitute the building blocks of these models [3]. These concepts, often regarded as mental representations, capture regularities in the world while abstracting away extraneous details, enabling flexible generalization to novel situations [4]. For example, the concept of sun can be instantly formed and deployed across diverse contexts: observing it rise or set in the sky, yearning for its warmth on a chilly winter day, or encountering someone who exudes positivity. Its role in the solar system can be analogized to the nucleus in an atom, enriching learning and understanding. The nature of concepts has long been a focus of inquiry across philosophy, cognitive science, neuroscience, and linguistics [5, 6, 7, 8, 9, 10, 11, 12]. These investigations have uncovered diverse properties that concepts need to satisfy, often framed by the long-standing divide between symbolism and connectionism. Symbolism emphasizes discrete, explicit symbols with structured and compositional properties, enabling abstract reasoning and recombination of ideas [13, 14]. In contrast, connectionism conceptualizes concepts as distributed, emergent patterns across networks, prioritizing continuity and gradedness, which excel in handling noisy inputs and learning from experience [15, 16]. Although there is growing consensus on the need to integrate the strengths of both paradigms to account for the complexity and richness of human concepts [17, 18], reconciling the competing demands remains a significant challenge.
@@
Recent advances in large language models (LLMs) within artificial intelligence (AI) have exhibited human-like behavior across various cognitive and linguistic tasks, from language generation [19, 20] to decision-making [21] and reasoning [22, 23, 24]. It has sparked intense interest and debate about whether these models are approaching human-like cognitive capacities based on concepts [25, 26, 27, 28, 29, 30]. Some argue that LLMs, trained exclusively on text data for next-token prediction, operate only on statistical associations between linguistic forms and lack concept-based understanding grounded in physical and social situations [31, 32, 29]. This argument is evidenced by their inconsistent performance, which often reveals vulnerabilities such as non-human-like errors and over-sensitivity to minor contextual variations [21, 33, 34]. Conversely, others contend that the performance of a system alone is insufficient to characterize its underlying competence [35], and the extent to which concepts should be grounded remains open [36, 37]. Instead, language may provide essential cues for people’s use of concepts and enable LLMs to approximate fundamental aspects of human cognition, where meaning arises from the relationships among concepts [38, 39]. Despite the conflicting views, there is broad consensus on the central role of concepts in human cognition. The core questions driving the debates are whether LLMs possess human-like conceptual representations and organization and, consequently, whether they can illuminate the nature of human concepts.
@@
To address these questions, we investigated the emergence of human-like conceptual representations from language prediction in LLMs. Our approach unfolded in three stages. First, we examined LLMs’ capacities to derive conceptual representations, focusing on the definitional and structured properties of human concepts. We leveraged LLMs’ in-context learning ability and guided them to infer concepts from descriptions with a few contextual demonstrations (Fig. 1). The models’ outputs offered a behavioral lens to evaluate the derived representations. We explored the organization of these representations by analyzing how their relational structures varied across different contextual demonstrations. Second, we assessed how well the LLM-derived representational structures align with psychological measures and investigated whether they capture rich, context-dependent human knowledge. These questions were tested by leveraging computations over the representations to predict human behavioral judgments. Finally, we mapped the LLM-derived conceptual representations to neural activity patterns in the human brain and analyzed their biological plausibility. Our experiments spanned thousands of naturalistic object concepts and beyond [40], providing a comprehensive analysis of LLMs’ conceptual representations. The findings reveal that language prediction can give rise to human-like conceptual representations and organization in LLMs. These representations integrated many valued aspects of previous accounts, combining the definitional and structural focus of symbolism with the continuity and graded nature of connectionist models. This work underscores the profound connection between language and human conceptual structures and demonstrates that LLM-derived conceptual representations offer a promising foundation for understanding human concepts.
@@ 2.1 Reverse dictionary as a conceptual probe
To guide LLMs through the process, we took advantage of their in-context learning ability and presented them with a small number of demonstrations in a reverse-dictionary format (“$[\textrm{Description}]\Rightarrow[\textrm{Word}]$”), followed by a query description (Fig. 1). We then compared model-generated completions to the intended term of the concept corresponding to the query description, thereby evaluating how well the models derived conceptual representations that aligned with human understanding. Formally, given $N$ pairs of descriptions and words as contextual demonstrations, an LLM $\mathcal{M}$ encodes the query description of a concept $c$ into a representation $\mathbf{h}_{c}=\textrm{encode}_{\mathcal{M}}\left(\textrm{description}\mid\textrm{context}_{N}\right)$, and generates a term based on it: $\textrm{predict}_{\mathcal{M}}\left(\mathbf{h}_{c}\right)$. The term was then compared to the name of the query concept shared by humans: $\textrm{name}_{\mathcal{H}}\left(c\right)$. We took $\mathbf{h}_{c}$ as the representation for concept $c$, which immediately precedes the following term that should be in semantic correspondence to the query description. Importantly, the demonstrations provided minimal and controllable context, and the query description served as a special case of general input for concept inference. We can analyze how models construct conceptual representations in response to different contextual cues by varying the number of demonstrations and the specific description-word pairings within them.
@@ 2.2 Deriving concepts from definitional descriptions
As a follow-up, we conducted a counterfactual analysis to investigate how LLMs accomplish this: Do they merely function as lookup tables, rigidly mapping descriptions to words, or do they contextually infer concepts based on their interrelationships? As shown in Fig. 2, when presented with only a single demonstration identical to the query description, the model replicated the context, even though the description was paired with a proxy symbol instead of the correct term. As additional correct demonstrations of other concepts were introduced, the model gradually shifted from replication to predicting the proper word for the query concept. This behavior persisted across various proxy symbols, suggesting that contextual cues about other concepts successfully guided the model in disregarding misleading information. When the correct demonstration for the query concept was provided, model performance slightly declined as more demonstrations were added, eventually dropping below the average in the standard generalization setting. This decline indicated that subtle conflicts among demonstrations could arise, with the influence of other concepts overshadowing that of the identical query concept. Collectively, these findings underscore the intricate interrelationships among concepts and their pivotal role in shaping model inference.
Figure 2: Performance of LLaMA3-70B on the reverse dictionary task measured through exact match accuracy. Across the four plots, the black lines denote the model’s performance when presented with $N$ demonstrations sampled from the training set and evaluated on an independent test set. The blue and red lines illustrate the model’s behavior when presented with one misleading demonstration—a description paired with a proxy label (as specified in the figure titles)—and $N-1$ correct demonstrations of other concepts. The blue line shows the frequency with which the model copies the misleading demonstration and reproduces the proxy label, while the red line indicates how often it generates the correct word for the query concept given contextual information from other concepts. Shaded areas denote 95% bootstrapped confidence intervals, calculated from 10,000 resamples over five independent runs.
@@ 2.3 Uncovering a context-independent conceptual structure
To uncover how concepts are represented and organized within LLMs, we looked into the representation spaces they constructed for concept inference and analyzed the interrelationships among the conceptual representations. We characterized the representation spaces formed under different contexts by their relational structure, captured through a (dis)similarity matrix, and measured pairwise alignment by correlating these matrices [43, 44] (Methods 4.3). For concepts in the THINGS database, the alignment between representation spaces gradually increased as the number of demonstrations rose from 1 to 24, with diminishing gains beyond this threshold (Fig. 3a). The alignment with the space formed by 120 demonstrations increased from $0.800$ (±$\pm 0.017$) with one demonstration to $0.970$ (±$\pm 0.003$) after 24 demonstrations. This alignment demonstrated a strong correlation with the model’s exact match accuracy on concept inference ($\rho=0.976$, $P<0.0001$, $95\%$ CI: 0.7300.7300.7300.7301.01.01.01.0) (Fig. 3b). These results suggest that LLMs are able to construct a coherent and context-independent relational structure, which is reflected in their concept inference capacity. Additional metrics corroborated this conclusion (Methods 4.3 and Supplementary Information SI 1.2), highlighting that intricate relationships among concepts are consistently preserved within LLMs’ representation spaces across varying contexts. This invariance supports the generalization of knowledge encoded in these relationships.
@@
Figure 3: LLMs converge toward a similar representational structure of concepts. a, Alignment correlation (RSA) between the LLM-derived conceptual representations across different contextual demonstrations. The LLM (LLaMA3-70B) is presented with $N$ demonstrations, repeated across five independent runs. b, LLM performance on the reverse dictionary task reflects alignment with the representations formed based on 120 demonstrations. Each point corresponds to the representations formed by the LLM based on $N$ demonstrations, with the x-axis indicating exact match accuracy and the y-axis denoting the alignment correlation. Error bars indicate $95\%$ confidence intervals, calculated using 10,000 bootstrap resamples from five independent runs. c, A t-SNE visualization of conceptual representations produced by different LLMs. Each point corresponds to an LLM, with representations formed based on 24 demonstrations. Distances between them are calculated as $1-\textrm{alignment}$, with results averaged over five independent runs. LLMs are plotted in proportion to their complexity and color-coded by their exact match accuracy on the reverse dictionary task. Better-performing models (blue) exhibit more aligned representations. d, LLMs’ concept inference capacity reflects alignment correlation with conceptual representations derived from LLaMA3-70B. The x-axis represents the average performance on the reverse dictionary task with 24 demonstrations, while the y-axis indicates the alignment correlation with LLaMA3-70B representations. Data points correspond to conceptual representations derived from different models and are color-coded by model complexity. Linear fits are shown as straight lines, with shaded areas representing $95\%$ confidence intervals derived from 10,000 bootstrap resamples. Overall, models with higher complexity align more closely with the reference. An exception is Qwen2-0.5B, which, despite exhibiting relatively high complexity, has a low alignment score. This may be due to constraints in model scale and the quality of training data.
@@ 2.4 Predicting various facets of human concept usage
Next, we investigated how well the conceptual representations and structures derived from LLMs align with various aspects of human concept usage. Specifically, we used these representations to predict human behavioral data across three key psychological phenomena: similarity judgments, categorization and gradient scales along various features. To evaluate their effectiveness, we compared the LLM-derived representations against traditional static word embeddings learned from word co-occurrence patterns [46].
@@
For similarity judgments, we compared LLM-derived conceptual representations with human similarity ratings for concept pairs from SimLex-999 [47], which has proven challenging for traditional word embeddings. We also complemented our analysis using human triplet odd-one-out judgments from the THINGS-data collection [40], which relied on image-based rather than text-based stimuli and introduced a contextual effect through the third concept. As shown in Fig. 4a, similarity scores derived from LLM representations strongly correlated with human ratings in SimLex-999, with the correlation improving as the number of contextual demonstrations increased ($\rho=0.776\pm 0.007$ with 72 demonstrations across five runs, $P<0.0001$). These representations significantly outperformed traditional word embeddings, which achieved a correlation of $\rho=0.464$ ($P<0.0001$). For THINGS triplets, we calculated pairwise similarities to identify the odd-one-out (Methods 4.5). The LLM prediction accuracy also improved with increasing contextual demonstrations, plateauing at $63.20\%$ (±$\pm 0.29\%$) after 48 demonstrations (Fig. 4b). This performance closely approached the noise ceiling estimated from individual behavior ($67.67\%\pm 1.08\%$) and substantially exceeded that of word embeddings ($48.10\%$). These findings suggest that, within proper context, the conceptual representations formed by LLMs effectively support the computation of similarities, a core property of human concepts.
@@
We then examined whether categories can be induced from the relative similarity between LLM-derived conceptual representations, using the human-labeled high-level categories from the THINGS database [42]. Applying a prototype-based categorization approach (Methods 4.6), we observed that LLM-derived representations consistently achieved high accuracy, reaching $92.25\%$ (±$\pm 0.15\%$) with only 24 demonstrations, significantly outperforming static word embeddings ($77.88\%$) (see Supplementary Information SI 1.4 and Fig. S10c). A t-SNE visualization of these representations revealed notable differences in their similarity structures (Fig. 4c–d). The LLM-derived conceptual representations formed distinct clusters corresponding to high-level categories, such as animals and food, while word embeddings exhibited less distinct category separation. Furthermore, LLM representations revealed broader distinctions, separating natural and animate concepts from man-made and inanimate ones. Among these groupings, human body parts clustered more closely with man-made objects than other animals do, while processed food appeared closer to natural and animate objects. These patterns align with previous findings on human mental representations [48] and are prominent in LLM representations, underscoring the meaningful correspondence between LLM-derived conceptual structures and human knowledge.
@@
Figure 4: Evaluation of alignment between LLM-derived conceptual representations and psychological measures of similarity. ab, Performance of LLaMA3-70B conceptual representations compared to static word embeddings (Word) in predicting human similarity judgments. a, Performance on human similarity ratings for concept pairs from SimLex-999, evaluated by Spearman’s rank correlation. b, Performance on human triplet odd-one-out similarity judgments in THINGS, measured by prediction accuracy. The noise ceiling reflects the upper bound of performance based on individual response consistency for the same triplet. Error bars and shaded regions represent 95% confidence intervals calculated from five independent runs. cd, t-SNE plots of LLM-derived conceptual representations (c) and static word embeddings (d). Data points are color-coded based on human-labeled categories from THINGS. Conceptual representations derived from LLaMA3-70B show a clear alignment with category structure, in contrast to the less distinct clustering observed with word embeddings.
@@
Finally, we investigated whether LLM conceptual representations could capture gradient scales of concepts along various features. For example, on a scale of 1 to 5 relative to other animals, cheetahs might rank a 5 for speed but fall closer to 3 for size. Using LLM-derived representations for such ratings, we predicted human behavioral data across 52 category-feature pairs (e.g., animals rated for speed) [49] (Methods 4.7). The results (Fig. 5) revealed strong correlations with human ratings ($\rho>0.5$, $P<0.001$, FDR controlled) for 48 out of 52 category-feature pairs, with a median correlation of $0.817$ ($95\%$ CI: 0.7770.7770.7770.7770.8510.8510.8510.851) across all pairs. When accounting for the split-half reliability of human ratings (median $\rho=0.954$), the median correlation reached $0.871$ ($95\%$ CI: 0.8410.8410.8410.8410.8910.8910.8910.891). Among the three category-feature pairs without statistically significant correlations (FDR $P>0.01$), two exhibited marginal significance (FDR $P<0.05$), while the weakest correlation appeared in the rating of clothing by age. As shown in Fig. 6, the LLM conceptual representations outperformed static word embeddings [49] across most category-feature pairs. This advantage remained robust after excluding concepts with extreme feature values from the correlation analysis (Supplementary Information SI 1.5), confirming that the success of LLMs is not driven by outlier items. Our findings demonstrate that LLM-derived conceptual representations advantageously handle intricate human knowledge across diverse object categories and features, highlighting their potential to advance our understanding of conceptual representations in the human mind.
@@
Figure 6: Comparison of LLM-derived conceptual representations and static word embeddings in predicting context-dependent human ratings. The x-axis represents correlations with human ratings based on static word embeddings, while the y-axis represents correlations based on conceptual representations derived from LLaMA3-70B. Colored points indicate significant correlations (Spearman’s rank correlation, FDR $P<0.01$). Error bars represent 95% confidence intervals estimated from 10,000 bootstrap resamples.